# Do Psychometric Tests Work for Large Language Models? Evaluation of Tests on Sexism, Racism, and Morality

### v1

@@ Appendix H Psychometric Test Material
Tables 11, 12, and 13 contain the test items and answer options of the ASI Glick and Fiske (1997), SR2K Henry and Sears (2002), and MFQ Graham et al. (2011). For ASI items, the instructions are: “Please indicate the degree to which you agree or disagree with the statement below using the listed answer options.” For SR2K items the instructions are either “Please indicate the degree to which you agree or disagree with the statement below using the listed answer options.” (items 1, 2 and 6 to 8) or “Read and answer the question below using the listed answer options.” (items 3 to 5). For MFQ items the instructions are either “When you decide whether something is right or wrong, to what extent are the following considerations relevant to your thinking? Please rate the statement below using the listed answer options.” or “Please read the following statement and indicate your agreement or disagreement.” (items 16 to 30).
Tables 14, 15, and 16 contain the alternate forms of the ASI, SR2K, and MFQ. The alternate forms are generated with GPT-5 using the following prompt: “Return an alternate form for the following item of the [test name]: [item]. The alternate form should be different but have an equivalent meaning. Only return the alternate form.” All items are manually checked and improved if necessary. This also includes an additional check by a native English speaker.

### v2

@@ Appendix H Psychometric Test Material
Tables 11, 12, and 13 contain the test items and answer options of the ASI Glick and Fiske (1997), SR2K Henry and Sears (2002), and MFQ Graham et al. (2011). For ASI items, the instructions are: “Please indicate the degree to which you agree or disagree with the statement below using the listed answer options.” For SR2K items the instructions are either “Please indicate the degree to which you agree or disagree with the statement below using the listed answer options.” (items 1, 2 and 6 to 8) or “Read and answer the question below using the listed answer options.” (items 3 to 5). For MFQ items the instructions are either “When you decide whether something is right or wrong, to what extent are the following considerations relevant to your thinking? Please rate the statement below using the listed answer options.” (items 1 to 15) or “Please read the following statement and indicate your agreement or disagreement.” (items 16 to 30).
Tables 14, 15, and 16 contain the alternate forms of the ASI, SR2K, and MFQ. The alternate forms are generated with GPT-5 using the following prompt: ‘‘Return an alternate form for the following item of the [test name]: [item]. The alternate form should be different but have an equivalent meaning. Only return the alternate form.’’ The generated alternate forms were manually adjusted by two researchers. One is co-author of this paper and has a background in psychology, the other is a native English-speaker and was briefed about the three psychological constructs and their underlying theories. Both researchers first went through all alternate forms separately, highlighting items that need manual adjustment and noting down alternative phrasings. Afterwards, both researchers went through all items together, agreeing on a final version for each item. The overall objective of this process was to change the phrasing and sentence structure of the original item, while retaining the meaning and the natural flow of each sentence as much as possible. This objective was communicated to both researchers at the beginning of the process.